Find the vertex , maximum , range and zeros of the formula f(x) = -x^2 + 2x +8
Thanks so much in advance:)

Answer by LinnW(1048)

f(x) = -x^2 + 2x +8
x value of vertex -b/2a = -2/2(-1) = 1
y value of vertex -1^2 +2*1 + 8 = 1 + 2 + 8 = 11
so the vertex is (1,11)
The maximum y value is 11 so the range is ( - infiniti , 11 )
To find the zeros
-x^2 + 2x + 8 = 0
multiply by -1
x^2 -2x -8 = 0
(x - 4)(x + 2) = 0
So x = 4 or x = -2